OpenStudy (anonymous):

a body is found at 8 am with a temp of 84, an hour later the coroner examines the body and the temp was 72. Assuming the body was 98 at death, and there was a constant temp of 35 degrees all night, what was the approx time of death?

OpenStudy (anonymous):

6:40am?

OpenStudy (anonymous):

how did you get that though? what equation did you use?

OpenStudy (anonymous):

I didn't use an equation.. Just kinda thought it out. If at 8am you checked the body and the temp was 84 degrees, then an hour later it was 72 degrees.. It dropped 12 degrees. So if he was 98 degrees when he died he would have lost the heat in his body at the same rate.. 12 degrees per hour. 98-12=86 degrees.. But at 8am he was 84. So it was more than an hour because he was colder than 86 degrees. So I used 6:40 because it seems reasonable. It makes so much more sense in my head. I'm sorry.

OpenStudy (anonymous):

lol know it's ok. It's really hard, it's like newtons law of cooling or something

OpenStudy (anonymous):

what I got, but I figure is wrong is 12=14e^1k but it didn't exactly compute cause it said he died 8 hours earlier

OpenStudy (anonymous):

Whoa! hahaha. Then definitely don't take my word for it. But for the record I change my guess to 6:50am, I thought about it more.

OpenStudy (anonymous):

Where does it say he died 8 hours earlier?

OpenStudy (anonymous):

well you get e^1k=14/12 from there you get 1k=ln(14/12) and that gives you .154 you use that for 26=98e^.154t and that comes out to -8.6

OpenStudy (anonymous):

but that was in comparison to a turkey cooking problem. This is opposite and it doesn't take the constant temp of 35 that night into the equation so I don't know what to do

OpenStudy (anonymous):

Ohhh.. That was the answer you got from the equation you used. Hmm. I don't know what to tell you. I probably would have done what I did if I were in your class too. I don't follow the math rules..lol. I'm a rebel. But my guess is probably wrong..lol. Best of luck. I have some work due at 12 tonight but if you still havent figured it out i'll look up some equations and try to help

OpenStudy (anonymous):

lol it's cool, have a nice night